# NUDT4MSTAR: A Large Dataset and Benchmark Towards Remote Sensing Object Recognition in the Wild

### v2

@@ Abstract
As an indispensable sensor for Remote sensing, Synthetic Aperture Radar (SAR) has a unique capability for all-day imaging. Nevertheless, in a data-driven era, the scarcity of large-scale datasets poses a significant bottleneck to advancing SAR automatic target recognition (ATR) technology. This paper introduces NUDT4MSTAR, a large-scale SAR dataset for remote sensing target recognition in the wild, including 40 vehicle target types and various imaging conditions across 5 realistic scenes. NUDT4MSTAR represents a significant leap forward in dataset scale, containing over 190,000 images—tenfold the size of its predecessors. We meticulously annotate each image with detailed target information and imaging conditions. Besides, data in both processed magnitude images and original complex formats are provided. Then, we construct a comprehensive benchmark consisting of 7 experiments with 15 recognition methods focusing on the stable and effective ATR issues. Besides, we conduct transfer learning experiments utilizing various models training on NUDT4MSTAR and apply them to three other target datasets, demonstrating its substantial potential for the broader field of ground objects ATR. Finally, we discuss this dataset’s application value and ATR’s significant challenges. To the best of our knowledge, this work marks the first-ever endeavor to create a large-scale dataset benchmark for fine-grained SAR recognition in the wild, featuring an extensive collection of exhaustively annotated vehicle images. We expect that the open source of NUDT4MSTAR will facilitate the development of SAR ATR and attract a wider community of researchers.
@@
Remote sensing, transformer, benchmark dataset, performance evaluation, synthetic aperture radar, automatic target recognition, image classification, object detection, foundation model, deep learning
@@ 1 Introduction
Figure 2: Motivation of our NUDT4MSTAR. Subfigure (a) depicts the most frequent keywords in 21,780 journal papers published in the remote sensing field (TGRS, J-STARS, GRSL, ISPRS P&R, and JAG) from 2020 to 2024. The size of each word is proportional to its frequency, highlighting that concepts such as synthetic aperture radar (SAR), image classification, and object detection [21, 10, 99] have garnered substantial attention. Subfigure (b) focuses on the number of publications related to SAR ATR over the past five years, a cross-area of the concepts highlighted in Subfigure (a). As the pioneering dataset for SAR target classification, MSTAR has long served as the predominant benchmark due to its unique data diversity and accumulated benchmarks. However, the lack of large-scale datasets has significantly limited the growth of this research field in recent years.
Remote sensing imagery interpretation is crucial for understanding our Earth comprehensively via diverse multi-modal sensors [64, 27, 111]. Recently, the advent of the big data era has propelled the evolution of pre-trained foundation models, which have been widely applied across various remote sensing downstream tasks involving different multi-modal data [30, 107, 87, 26, 3, 58]. Nevertheless, not all modal tasks in remote sensing possess a large number of target samples to exploit the potential of big data and foundation models.
Synthetic aperture radar (SAR) [78, 36], an active imaging sensor, acquires the ground scattering characteristics by emitting the electromagnetic microwave. As shown in Fig 2 (a), it has emerged as an indispensable tool for information acquisition in Remote sensing and Earth observation [23, 83, 21], as its resolution remains unaffected by the observation distance or lighting conditions. SAR automatic target recognition (ATR), which autonomously detects and classifies objects of interest (e.g., , vehicles, ships, aircraft, and buildings) within imagery, is a critical increasing frontier in SAR interpretation, as illustrated in Fig. 2 (b). It has various civil and military applications, such as urban management, disaster assessment, emergency rescue, and global surveillance [10, 54, 28, 108]. Can humans perceive and recognize objects on the Earth beyond the limitations of optical vision using microwave vision [21, 36, 49]? Since the 1970s, researchers have been dedicated to developing a generalized stable SAR ATR system. However, recognizing objects within SAR images has encountered significant challenges in a big data era, especially the scarcity of large-scale target samples.
Special characteristics. SAR imagery has special electromagnetic scattering attributions but is devoid of common features such as color and texture typically present in human vision. The scattering properties of targets are highly influenced by a multitude of factors, including the imaging angle, resolution, polarization, background, and the geometry structure and material of targets themselves [73, 44, 69]. Additionally, SAR images suffer quality issues, such as speckle noise, geometric distortion, and image Defocus. These factors make it difficult to recognize objects in SAR, a unique recognition system independent of human vision.
Insufficient samples. Acquiring SAR images encompassing many categories under diverse operating conditions presents a significant challenge, and annotating SAR images is considerably more costly and labor-intensive than RGB images. Three decades have elapsed since the advent of the first SAR classification dataset in the 1990s, the Moving and Stationary Target Acquisition and Recognition (MSTAR) [2]. As shown in Fig. 3, the largest publicly accessible SAR target classification dataset remains limited to a small set of fewer than 30,000 images. Besides, recent classification datasets (OpenSARShip [34] and FUSAR-Ship [33]) have severe sample imbalance, ranging from a few to thousands of samples per class. A substantial amount of work remains to be undertaken for SAR target datasets.
Nonstandard benchmarks. A standardized dataset and experimental settings serve as the north star for developing advanced SAR ATR techniques. However, the mainstream MSTAR dataset in Fig. 2 (b) has dozens of distinct experimental settings, and existing algorithms have nearly reached near-saturation performance on this ideal benchmark [44]. Other classification datasets, such as OpenSARShip and FUSAR-Ship, also exhibit different experimental settings and dataset splitting. Moreover, the limiting access to non-open-source dataset settings and methods further intensifies this challenge.
Researchers have struggled to construct many SAR target datasets to study generalized stable SAR ATR algorithms. Over the past few years, there has been a remarkable surge in remote sensing datasets across various modalities, including visible light [5, 56], SAR [57, 93], infrared [98], and multimodal [65]. Recent SAR classification datasets in Table II have significantly improved data diversity. In particular, SAR detection datasets [57, 93] with 100,000 images have emerged. Yet, regarding SAR target recognition, no dataset has emerged that can completely surpass the MSTAR benchmark dataset, released in the 1990s, regarding the diversity of target types, imaging conditions, and data formats it offers. Notably, the lack of open-source availability, inadequate samples, long-tailed distributions, unabundant collection conditions, and independent experimental settings are still all issues that warrant attention. We suggest that a large-scale fine-grained SAR target dataset and benchmarks to investigate the electromagnetic scattering properties of targets across diverse conditions represent a critical and enduring endeavor.
@@
TABLE I: MSTAR v.s. NUDT4MSTAR. We can see the potential of the NUDT4MSTAR dataset as a new benchmark, surpassing MSTAR in terms of data diversity and challenges. AFRL: Air Force Research Laboratory. NUDT: National University of Defense Technology. # Types.: Number of object types. # Samples: Number of samples.
Dataset MSTAR NUDT4MSTAR
@@
As a first step towards building a large-scale SAR target dataset, we introduce NUDT4MSTAR [^1] The name NUDT4MSTAR comes from the abbreviation of the College of Electronic Science and Technology, the National University of Defense Technology, as Department IV. This dataset focuses on the finegrained single target recognition problem in SAR ATR, i.e., the image classification task, but we provide a horizontal bounding box label for the object detection task. NUDT4MSTAR: https://github.com/waterdisappear/NUDT4MSTAR., a fine-grained dataset for vehicle target recognition as a new breakthrough to the previous benchmark dataset MSTAR in Table. I. In particular, NUDT4MSTAR collects 194,324 object images in the wild from 40 target types, 5 scenes, and various imaging conditions with detailed annotation. It is the largest public SAR vehicle recognition dataset, even 10 times the size of any similar vehicle dataset previously available, as shown in Table. II. Then, to facilitate innovation and comparisons, we build a well-designed benchmark with 7 experimental settings and 15 recognition methods. Besides, the transfer learning results of various models trained on this dataset to other datasets show that our fine-grained recognition task benefits the feature learning of different ground targets. Finally, We provide detailed analyses and discussions to explore important issues for future research.
Our work shows the huge challenges and potential in SAR recognition and provides the benchmark for stable fine-grained ATR algorithms. The dataset and benchmark are available on our GitHub. We believe that our new dataset, with comprehensive imaging conditions and experiment benchmark, will further promote the promise and reproducible research in SAR ATR. The main contributions of this paper are three points:
1. We build a large dataset for remote sensing object recognition in the Wild. NUDT4MSTAR includes 40 distinct vehicle target types and over 194,324 target samples across 5 realistic scenes and various imaging conditions. It provides a systematic study of the effects of different conditions on SAR target characteristics and promotes the development of new ATR research topics.
2. We provide a comprehensive benchmark with 7 experimental settings and 15 recognition methods, covering both classification and detection recognition tasks. This benchmark offers a standardized experimental platform for researchers to compare the performance of different algorithms, facilitating innovation and development in SAR ATR technology.
3. We demonstrate the NUDT4MSTAR dataset’s potential for foundation models. A transfer learning experiment based on various models pre-trained on this dataset shows that it is not only suitable for vehicle target recognition but also has the potential to support research in other related objects, promoting the application of SAR ATR in multiple situations.
@@
The remainder of this paper is organized as follows. Sec. 2 discusses related work in SAR target classification datasets and algorithms. Sec. 3 introduces our dataset NUDT4MSTAR. Secs. 4 conduct extensive experiments to build the benchmark. Sec. 5 concludes the paper and discusses future work.
@@ 2.1 Datasets for SAR target classification
Datasets are the cornerstone of developing new techniques for SAR ATR, especially in a deep learning-based era. In the past decades, the MSTAR [2] dataset has played a key role in SAR target classification tasks. However, current deep learning-based methods are saturated on this dataset benchmark setting. MSTAR’s insufficient data volume also suffers from data biases such as background correlation [44, 52]. Therefore, many datasets have been built to advance the development of this field, especially in recent years. In order to provide a comprehensive analysis, we discuss SAR target classification datasets that focus on individual target characteristics from the 1990s to the 2020s, which are detailed listed in Table II and Fig. 3.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Vehicle dataset. Vehicle datasets usually include various types of fine-grained vehicles and imaging conditions to study the robustness classification under different distributions in the training and test sets. We summarize the target classes and types of these SAR vehicle datasets in Fig. 5. MSTAR was released by the Defense Advanced Research Projects Agency (DARPA) as the first public dataset for SAR target classification. It contains SAR image slices of 10 military vehicles and a reference target under different imaging angles with spotlight SAR. However, MSTAR has a background correlation problem and ideal collection conditions. Compared to MSTAR, the QinetiQ [74] dataset provided a non-idealized acquisition condition with a non-centered target location and stronger background clutter. CV Domes [18] and Gotcha [17] constructed simulation and measurement samples of civilian vehicles. These two datasets are in the raw echo data format and can be used for SAR imaging research, but they are not as convenient for researchers to use as MSTAR. However, these two civilian vehicle datasets are not rich in target categories and acquisition conditions. SARSIM [66] and SAMPLE [47] are two simulation datasets used to study transfer learning. In particular, SARSIM uses 14 CAD models, and SAMPLE has simulated and measured sample pairs based on partially MSTAR’s public vehicle. The simulation data have discrepancies with the measured data in background clutter, target scattering characteristics, and speckle noise. Although MSTAR includes 3 grass locations, most public images are from the same location. MGTD [6] collects the training and test sets in two different laboratory backgrounds. Another issue is that the depression angles of MSTAR public images are mostly in 15° and 17°, and some categories of images are missing at other depression angles, such as 30° and 45°. Hence, NUAAminiSAR [112] collects images of military targets at different depression angles by circular SAR in a national defense park, including 8 military vehicles and an airplane. As shown above, no vehicle datasets are sufficiently rich in target types, scene settings, and sensor conditions.
@@
Ship dataset. SAR ship datasets have emerged with the development of satellite SAR technology. The difference with the vehicle dataset is that the ship classification dataset usually uses a satellite platform, and the imaging angles are limited. Ship datasets can be constructed and labeled with global ship tracking systems, but label noise is a difficult problem. OpenSARShip [34], HR4S [4], and FUSAR-Ship [33] advance the development of ship classification methods. These datasets are derived from different satellites and face difficult harbor backgrounds and class imbalance challenges. OpenSARShip collects medium-resolution images of 16 classes in harbors with Sentinel-1. HR4S constructs high-resolution ship images in ports and coastal areas with GF-3 and RADARSAT-2. FUSAR-Ship contains high-resolution images under various sea backgrounds with 15 ship main classes and 98 subclasses based on GF-3. However, OpenSARShip and FUSAR-Ship have a serious class imbalance and long-tailed distribution challenges. Researchers construct various experimental settings and use some target categories for the classification based on these two datasets.
@@
Aircraft dataset. Aircraft datasets have emerged in recent years, presenting unique challenges in remote sensing. Due to their streamlined surfaces and less obvious scattering characteristics, aircraft are more difficult to classify than vehicle and ship targets. Multiangle Aircraft Target Dataset (MATD) [89] collects SAR images of two aircraft at different imaging angles with a drone strimap SAR equipment. SAR-ACD [79] contains 2 categories and 6 types of civil aircraft in 3 airports.
@@
Object detection dataset. SAR has many object detection datasets besides the classification datasets used for individual targets. These datasets are used for detecting multiple targets and have large image sizes. They contain more background interference and intensive targeting challenges. SAR object detection datasets similar to SARDet-100K [57] and FAIR-CSAR [93] with a volume of about 100,000 images are mainly satellite-based datasets, with a resolution of mostly greater than 1 m. This resolution makes it hard to detect and classify vehicle targets. Their non-cooperation objectives make it difficult to label the target types and accurately control the characteristics of the target under different angles and scenes. Therefore, our dataset provides a more fine-grained SAR vehicle characterization dataset under target class and imaging conditions.
Due to the high collection and annotation cost, collecting and publishing target samples under different imaging conditions is very difficult and time-consuming. Even after close to thirty years since its inception, MSTAR remains the most used classification dataset. The evolution of SAR ATR necessitates a new large-scale dataset and benchmark in the new era. After the above investigation, we suggest that a good SAR ATR dataset should pay attention to the following key points.

### v3

# ATRNet-STAR: A Large Dataset and Benchmark Towards Remote Sensing Object Recognition in the Wild

@@ Abstract
The absence of publicly available, large-scale, high-quality datasets for Synthetic Aperture Radar Automatic Target Recognition (SAR ATR) has significantly hindered the application of rapidly advancing deep learning techniques, which hold huge potential to unlock new capabilities in this field. This is primarily because collecting large volumes of diverse target samples from SAR images is prohibitively expensive, largely due to privacy concerns, the characteristics of microwave radar imagery perception, and the need for specialized expertise in data annotation. Throughout the history of SAR ATR research, there have been only a number of small datasets, mainly including targets like ships, airplanes, buildings, etc. There is only one vehicle dataset MSTAR collected in the 1990s, which has been a valuable source for SAR ATR. To fill this gap, this paper introduces a large-scale, new dataset named ATRNet-STAR with 40 different vehicle categories collected under various realistic imaging conditions and scenes. It marks a substantial advancement in dataset scale and diversity, comprising over 190,000 well-annotated samples—10×10$\times$10 × larger than its predecessor, the famous MSTAR. Building such a large dataset is a challenging task, and the data collection scheme will be detailed. Secondly, we illustrate the value of ATRNet-STAR via extensively evaluating the performance of 15 representative methods with 7 different experimental settings on challenging classification and detection benchmarks derived from the dataset. Finally, based on our extensive experiments, we identify valuable insights for SAR ATR and discuss potential future research directions in this field. We hope that the scale, diversity, and benchmark of ATRNet-STAR can significantly facilitate the advancement of SAR ATR.
@@
Remote sensing, transformer, datasets and benchmarks, performance evaluation, synthetic aperture radar, automatic target recognition, image classification, object detection, foundation model
@@ 1 Introduction
Figure 2: Motivation of our ATRNet-STAR. Subfigure (a) depicts the most frequent keywords in 21,780 journal papers published in the remote sensing field (TGRS, JSTARS, GRSL, ISPRS Journal, and JAG) from 2020 to 2024. The size of each word is proportional to its frequency, highlighting that concepts such as synthetic aperture radar (SAR), image classification, and object detection have garnered substantial attention. Subfigure (b) focuses on the number of publications related to SAR ATR over the past five years, a cross-area of the concepts highlighted in Subfigure (a). As the pioneering dataset for SAR target classification, MSTAR has long served as the predominant benchmark due to its unique data diversity and accumulated benchmarks. However, the lack of large-scale datasets has significantly limited the growth of this research field in recent years.
Synthetic Aperture Radar (SAR) imaging is capable of generating high resolution imagery irrespective of lighting conditions and weather [1, 2, 3, 4], and has become an indispensable tool for Earth Observation [5, 2, 6]. SAR Remote Sensing (RS) data enables analysis and recognition of objects and scenes [2, 3, 7], which has become a valuable complement to other RS imagery [8, 9, 10]. Consequently, as a fundamental and challenging field in RS image analysis, SAR automatic target recognition (ATR) [11, 12], which autonomously detects and classifies objects of interest (e.g., , vehicles, ships, aircraft, and buildings), has become an active research area for several decades (Fig. 2 (a)). SAR ATR has a wide range of civilian and military applications, including military reconnaissance, urban management, disaster assessment, emergency rescue, and global surveillance [13, 14, 15, 16]. Despite the remarkable achievement achieved over the past several decades in the field of SAR ATR, the accurate, robust, and efficient recognition of any target in an open world remains unresolved [11, 7].
Need for ATRNet The advent of the big data era has propelled the evolution of RS pretraining foundation models [17, 18, 19, 20] where large-scale pre-training enables efficient cross-task adaptation with minimal finetuning. However, the scarcity of large-scale standardized datasets restricts the development of generalizable data-driven SAR ATR methods, compared to the success of foundation models in other RS sensors [21, 22, 23]. The data sensitivity, acquisition costs, annotation difficulty, and complexity of SAR imaging hinder the establishment of open large-scale data ecosystems: More than 50% of SAR ATR studies still rely on the 1990s-era Moving and Stationary Target Acquisition and Recognition (MSTAR) [24] dataset (Fig. 2 (b)). In addition, non-standardized evaluation protocols across existing benchmarks (e.g., MSTAR, OpenSARShip [25], and FUSAR-Ship [26]) impede objective algorithm comparison. Therefore, building a large SAR ATR dataset and benchmark is necessary to unlock new model capabilities in this field. We aim for ATRNet—a massive, diversified SAR target dataset benchmark for modeling and recognizing target characteristics.
Need for ATRNet-STAR. Researchers have struggled to construct many SAR target datasets to study SAR ATR algorithms. Many SAR classification datasets in Table II have significantly improved data diversity. In particular, new SAR detection datasets [27, 28] have emerged with 100,000 images. However, our previous research [29, 30] on SAR foundation models revealed that collecting publicly available datasets yields fewer than 300,000 target samples, with severe sample imbalance: mainly ship detection datasets. As the inaugural phase of ATRNet, we focus on vehicle targets due to: (1) Recent datasets predominantly are targets with spaceborne SAR and constrained imaging geometries, whereas airborne platforms offer higher resolution (0.1-0.3m) and richer imaging conditions. (2) Vehicle recognition has driven SAR ATR research for three decades since the building of MSTAR, establishing mature research systems. (3) The 1990s-era MSTAR dataset, despite its seminal role, suffers from idealized acquisition conditions and saturated performance (near 99% accuracy), failing to reflect real-world complexities and support innovation in the 2020s.
@@
TABLE I: MSTAR v.s. ATRNet-STAR. We can see the potential of the ATRNet-STAR dataset as a new benchmark, surpassing MSTAR in terms of data diversity. AFRL: Air Force Research Laboratory. NUDT: National University of Defense Technology. # Types.: Number of object types. # Samples: Number of samples.
Dataset MSTAR ATRNet-STAR
@@
Our Objective and Solution. As a first step towards building a large-scale SAR target dataset, we introduce ATRNet-STAR [^1] STAR: Stationary Target Acquisition and Recognition. This dataset focuses on the fine-grained single target recognition problem in SAR ATR, i.e., the image classification task, but we provide a horizontal bounding box label for the object detection task in different condictions., a fine-grained dataset for vehicle target recognition as a new breakthrough to the previous benchmark dataset MSTAR in Table. I. We completed program designation, data acquisition and processing, and benchmark construction with nearly two years of effort and considerable resources. In particular, ATRNet-STAR collects 194,324 object images from 40 target types, 5 scenes, and various imaging conditions with detailed annotation. It is the largest public SAR vehicle recognition dataset, even 10 times the size of any similar vehicle dataset previously available, as shown in Table. II. Its sufficient target samples can support multiple aspects of generation, detection, and classification research. To facilitate innovation and comparisons, we built a well-designed classification and detection benchmark, including 15 representative methods based on 7 experimental settings for robust ATR, few-shot, and transfer learning. Our results show that SAR ATR under complex conditions is still very difficult, and representative methods that performed well in previous datasets can not solve some of our difficult experimental conditions, such as image angle and scene variations. We also find the effectiveness of the pre-trained foundation model compared to other methods, which illustrates the importance of a large-scale dataset. Besides, the transfer learning results of various models pre-trained on this dataset to other datasets show that our dataset benefits the feature learning of different ground targets. Finally, We provide detailed analyses and discussions to explore important issues for future research. We believe that our new dataset, with comprehensive imaging conditions and experiment benchmark, will further promote the promise and reproducible research in SAR ATR.
Our Contributions. The main contributions of this paper are listed below.
Dataset and Benchmarks. To our best knowledge, the ATRNet-STAR is the largest dataset towards SAR vehicle recognition in the wild. Building ATRNet-STAR is a challenging task, and has taken nearly two years. We derive benchmark datasets from ATRNet-STAR to facilitate the performance evaluation of target classification and detection methods.
Code library. To the best of our knowledge, we are the first to integrate publicly available and representative methods into a single code library to facilitate reproducible and large-scale performance evaluations in this field.
Performance Evaluation. We review most major approaches and provide a comparative study of 15 representative methods with 7 experimental settings on classification and detection benchmarks derived from ATRNet-STAR, offering official baseline references when developing new approaches.
Insights and Future Directions. By analyzing the quantitative results, we provide insightful discussion that can guide the development of SAR ATR algorithms and provide potential future research directions in this field.
@@
This work mainly focuses on the introduction, performance evaluation, and potential value of ATRNet-STAR. The full annotated dataset, code library, and reported results are available at https://github.com/waterdisappear/ATRNet-STAR after acceptance. The remainder of this paper is organized as follows. Sec. 2 discusses related work in SAR target classification datasets and algorithms. Sec. 3 introduces our dataset ATRNet-STAR. Secs. 4 conduct extensive experiments to build the benchmark. Sec. 5 concludes the paper and discusses future work.
@@ 2.1 Datasets for SAR target classification
Datasets are the cornerstone of the development of new ATR techniques, especially in an era based on deep learning. In recent years, there has been notable progress in RS datasets across various modalities [31], including visible light [32, 33, 34, 35, 36], SAR [27, 28], infrared [37], and multimodal [38, 39]. However, not all sensors possess sufficient target samples to fully leverage the potential of big data and foundation models. In the past decades, the MSTAR [24] dataset has played a pivotal role in SAR target classification tasks. However, current deep learning-based methods are saturated on this dataset benchmark setting. MSTAR’s insufficient samples also suffers from data biases such as background correlation [11, 40, 41]. Therefore, many datasets been developed to promote the advancement of this field, especially in recent years. We analyze SAR target classification datasets that focus on individual target characteristics from the 1990s to the 2020s, which are detailed listed in Table II and Fig. 3.
@@
MSTAR [24] 1995 $\checkmark$ airborne X single 8 10 0.3 128$\sim$193 14,557 Most cited vehicle dataset
QinetiQ [42] 2004 - airborne X quad - 9 0.3 100$\sim$150 4,006 Non-ideal background conditions
@@
OpenSARShip [25] 2018 $\checkmark$ satellite C dual 16 - 2.7$\sim$22 9$\sim$445 26,679 Sentinel-1 ship dataset
HR4S [46] 2019 - satellite C quad 21 - 3$\sim$25 - 1,962 GF-3 and RADARSAT-2 ship
@@
FUSAR-Ship [26] 2020 $\checkmark$ satellite C dual 98 - 1.1$\sim$1.7 512 5,243 High-resolution GF-3 ship
@@
SAR-ACD [52] 2022 $\checkmark$ satellite C single 2 6 1 32$\sim$200 3,032 GF-3 aircraft dataset
ATRNet-STAR 2024 $\checkmark$ airborne X, Ku quad 21 40 0.12 $\sim$ 0.15 128 194,324 A large-scale fine-grained dataset
@@
Vehicle datasets. Vehicle datasets typically include a wide variety of fine-grained vehicle types and diverse imaging conditions, designed to evaluate robust classification performance under varying training and test set. These SAR vehicle datasets are summarized in Fig. 5, which provides an overview of the target categories. MSTAR was released by the Defense Advanced Research Projects Agency (DARPA) as the first public dataset for SAR target classification. It contains SAR image slices of 10 military vehicles and a reference target under different imaging angles with spotlight SAR. However, MSTAR has a background correlation problem and ideal collection conditions. Compared to MSTAR, the QinetiQ [42] dataset provided a non-idealized acquisition condition with a non-centered target location and stronger background clutter. CV Domes [43] and Gotcha [44] constructed simulation and measurement samples of civilian vehicles. These two datasets are in the raw echo data format and can be used for SAR imaging research, but they are not as convenient for researchers to use as MSTAR. Besides, these two civilian vehicle datasets are not rich in target classes and acquisition conditions. SARSIM [45] and SAMPLE [48] are two simulation datasets designed to study transfer learning. In particular, SARSIM uses 14 CAD models, and SAMPLE has simulated and measured sample pairs based on partially MSTAR’s public vehicle. The simulation data have discrepancies with the measured data in background clutter, target scattering characteristics, and speckle noise. Although MSTAR includes 3 grass locations, most public images are from the same location. MGTD [47] collects the training and test sets in two different laboratory backgrounds. Another issue is that the depression angles of MSTAR public images are mostly in 15 and 17, with some classes missing at other depression angles. Hence, NUAAminiSAR [50] collects images of military targets at different depression angles by circular SAR in a national defense park, including 8 military vehicles and an airplane. As shown above, existing vehicle datasets lack sufficient diversity in target types, scene settings, and sensor conditions.
@@
Ship datasets. SAR ship datasets have emerged with the development of satellite SAR technology. The difference with the vehicle dataset is that the ship classification dataset usually uses a satellite platform, and the imaging angles are limited. Ship datasets can be constructed and annotated using global ship tracking systems, but label noise remains a significant challenge. OpenSARShip [25], HR4S [46], and FUSAR-Ship [26] advance the development of ship classification methods. These datasets are derived from different satellites and face difficult harbor backgrounds and class imbalance challenges. OpenSARShip collects medium-resolution images of 16 classes in harbors with Sentinel-1. HR4S constructs high-resolution ship images in ports and coastal areas with GF-3 and RADARSAT-2. FUSAR-Ship contains high-resolution images under various sea backgrounds with 15 ship main classes and 98 subclasses based on GF-3. However, OpenSARShip and FUSAR-Ship have a serious class imbalance and long-tailed distribution challenges. Researchers construct various experimental settings and use some target categories for the classification based on these two datasets.
@@
Aircraft datasets. Aircraft datasets have emerged in recent years, presenting unique challenges in RS. Due to their streamlined surfaces and less obvious scattering characteristics, aircraft are more difficult to classify than vehicle and ship targets. Multiangle Aircraft Target Dataset (MATD) [51] collects SAR images of two aircraft at different imaging angles with a drone strimap SAR equipment. SAR-ACD [52] contains 2 classes and 6 types of civil aircraft in 3 airports.
@@
Object detection dataset. Beyond the classification datasets focused on individual targets, numerous SAR object detection datasets are designed to detect multiple targets. These datasets typically feature large image sizes, significant background interference, and complex targeting challenges. Examples include SARDet-100K [27] and FAIR-CSAR [28], mainly satellite-based datasets containing approximately 100,000 images with resolutions exceeding 1 meter. This resolution makes it hard to detect and classify vehicle targets. Their non-cooperation objectives make it difficult to label the target types and accurately control the characteristics of the target under different angles and scenes. Therefore, our dataset provides a more fine-grained SAR vehicle characterization dataset under target class and imaging conditions.
Collecting target samples under different imaging conditions remains a challenging and resource-intensive endeavor. After thirty years, MSTAR remains the most used classification dataset. The evolution of SAR ATR necessitates a new large-scale dataset and benchmark in the new era. After the above investigation, we suggest that a good SAR ATR dataset should pay attention to the following key points.